ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: 94 Plümer, Judith, University of Osnabrück, Germany Mathematical Metadata Publishing an article in a mathematical journal takes quite some time. That is the reason why mathematicians make their scientific work available on the internet in form of preprints. Resource discovery of mathematical preprints should be faciliated. What one wants is high quality retrieval for high quality material. We show how the use of MetaData allows for a suitable search engine. Creation of MetaData can be done by authors or providers of preprint servers using freely available forminterfaces. In short we will demonstrate a prototype of MPRESS as the EMS will name the service for retrieval of distributed mathematical preprints in Europe. Reference: http://www.mathematik.uni-osnabrueck.de/harvest/brokers/FraGer/.